graph the function
y=[-3x]-1

Answer by haileytucki(390)

y=[-3x]-1
To find the x-intercept, substitute in 0 for y and solve for x.
(0)=[-3x]-1
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
[-3x]-1=(0)
Remove the parentheses around the expression -3x.
-3x-1=(0)
Remove the parentheses around the expression 0.
-3x-1=0
Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides.
-3x=1
Divide each term in the equation by -3.
-(3x)/(-3)=(1)/(-3)
Simplify the left-hand side of the equation by canceling the common factors.
x=(1)/(-3)
Simplify the right-hand side of the equation by simplifying each term.
x=-(1)/(3)
To find the y-intercept, substitute in 0 for x and solve for y.
y=[-3(0)]-1
Multiply -3 by each term inside the parentheses.
y=0-1
Solve the equation.
y=-1
These are the x and y intercepts of the equation y=[-3x]-1.
x=-(1)/(3), y=-1